Find the area under the normal curve in each of the following cases. between z = 0 and z = 1.63 between z= 0 and z = 1.78 between z = -2.46 and z = 1.55 to the right of z = 2.35 to right of z = -1.31 to the left of z = 0.35 to the left of z - 1.85 to the right of z = -0.95

1
Expert's answer
2022-04-11T18:10:19-0400

between z = 0 and z = 1.63

0.9484-0.5=0.4484

between z= 0 and z = 1.78

0.9625-0.5=0.4625

between z = -2.46 and z = 1.55

0.9394-0.0069=0.9325

to the right of z = 2.35

1-0.9906=0.0094

to right of z = -1.31

1-0.0951=0.9049

to the left of z = 0.35

0.6368

to the left of z - 1.85

0.0322

to the right of z = -0.95

1-0.1711=0.8289